# GPU Volume Rendering with Hierarchical Compression Using VDB

### v1

@@ 1 Introduction
In this paper we propose to hierarchically encode—and by doing so compress—the volume data, allowing for per-thread decompression in a GPU compute kernel. We use the sparse voxel representation (VDB) for that. VDBs, so named because they represent volumetric grids that share similarities with B+trees [Mus13], are popular in production rendering and industry-strength libraries exist that are optimized for construction and sampling on NVIDIA GPUs. As such, VDBs are often used for data sets like the one shown in GPU Volume Rendering with Hierarchical Compression Using VDB.
@@ 3.1 Fixed-Rate Compression Algorithm
Algorithm˜1 provides a high-level overview. The input consists of the volume itself, its spatial extent represented by $W$, $H$, and $D$ (in voxels), and the user-provided quality value. We define point sampling routines on the volume itself—without loss of generality, we only sample the volume at exact (integer) voxel positions though (volume.value); as well as routines to activate and set individual voxels on the VDB (Activate).
1:function Compress(volumevolume, $W$, $H$, $D$, qualityquality)
@@
3:  $I=$ argmax(histhist)
4:  $B=volume.valueAt(I)$ $\triangleright$ background value
@@
6:  $brickSize=$ int3int3(252**5)
7:  $numBricks=$ int3int3(W,H,D/brickSize{W,H,D}/brickSize)
@@
11:   lo=brickIDbrickSizelo=brickID*brickSize
12:   hi=(brickID+1)brickSizehi=(brickID+1)*brickSize
@@
16:  SortBy(brickRefsbrickRefs, valueRangesvalueRanges, SimilarityFunc($B$))
@@
19:  bricksToActivate=numBricksqualitybricksToActivate=numBricks*quality
@@
22:     Activate(VoxelVoxel, $volume.value(Voxel)$)
@@
In the second phase we form bricks of size $32\times 32\times 32$ to cover the original volume, strategically choosing this size to cover the leaf node level of the VDB. We compute value ranges for each brick using the original volume.
@@
$f1$ and $f2$ compute the closest and the farthest distance from the scalar to the range given by $[lo:hi]$, respectively, while $f3$ computes the distance to the median of the range. We refer to the metrics as the closest, farthest and median point-in-range metrics and evaluate their impact on compression in Section˜5.
@@ 3.2.1 Compression
We use OpenVDB to implement the fixed-rate compression algorithm. We start from an empty openvdb::FloatGrid. In OpenVDB, grids associate trees, the internal representation of the VDB in memory, with transforms (e.g., voxel to object space). This internal representation is a Tree4<float, 5, 4, 3>, i.e., the tree has four levels using the $\{3,4,5\}$ layout as described above. The tree provides a function addTile(level,Coord(),value) to set and activate cells on each tree level.
@@
Regarding compression performance we note a $1:1$ relationship between activated voxels and final output size. We also note that activating a whole tile (addTile(), see above) on a level higher than leaf node level 0 results in OpenVDB just setting all the voxels in that tile to the same value, so there is no additional compression to gain from using hierarchical or level-of-detail encoding here.
@@ 4 ANARI Renderer
Given our conversion algorithm plus texture sampling routines on the GPU we can implement a renderer. We chose to give the renderer an ANARI API. ANARI support has recently been added to ParaView [WZA24] so our renderer can be evaluated using standard visualization tools and be compared with other renderers that also use the ANARI standard (cf. Fig.˜1).
@@ 5 Evaluation
We evaluate the quality and performance of our compressed representation for direct volume rendering. For that we use the structured-regular volume data sets from Table˜1. The aneurism data set is particularly sparse. The WDAS cloud data set by Walt Disney Animation Studios [Wal18] is originally available as a VDB and was resampled to a structured-regular representation retaining the original 2K resolution. As we also have access to the original VDB, this gives us the possibility to compare back-and-forth conversion using our algorithm. The airplane data set was simulated with OpenFOAM; we use the Q-criterion vorticity field which is usually sparse compared to other variables. The data was voxelized using the MESIO library [MvJB22], as detailed in [FMB25]. We have access to two versions of that data set: one that represents the vorticity, and a filtered version of that field where $Q\in[0:500K]$, retaining only 4.4 % of the original voxels. We also use a high-resolution, non-cosmological isolated galaxy data set for the evaluation. We utilize an example from [WS23], designed to resemble a Milky Way-type galaxy. The simulation incorporates magnetohydrodynamics (MHD) along with standard subgrid physics commonly employed in galaxy formation studies, including gas dynamics with a magnetic field. The simulation was produces using ChaNga [JGM08] and subsequently converted to a floating-point volume on a structured-regular grid. To simplify the comparison we assume voxels to be represented with 32-bit floating-point values; we convert aneurism and heptane to that format whose voxels use 8-bit values.
@@ 5.1 Compression Rate
We first evaluate if, by using our fixed-rate encoding, the target compression rate is achieved and compress the data set with different ratios from 1:8 to 1:1 (the latter meaning the targeted reduction in size is zero percent). In the latter case the algorithm is guaranteed to be lossless but will only activate voxels that are not considered empty, so the actual size will be lower than the original one. To assess this we report the actual size in memory of the resulting NanoVDBs in Table˜2.
@@
We are also interested in the worst case we have to expect when losslessly compressing a volume that is not sparse. This is an extreme case that we simulate by generating synthetic volume data sets. We generate noise textures of different sizes to fill the structured-regular grids with voxels so that no local neighborhood is considered empty by our algorithm. We report results for this experiment in Table˜3. For large-scale data we observe that the overhead does not exceed 10 % of the original size. We also found, although without a formal evaluation, that there is no correlation between sparseness and render performance and that NanoVDB performs equally well on data that is dense. This suggests that VDBs are a viable replacement even for dense textures.
@@ 5.2 Compression Quality
Table 4: Compression rate statistics. We report mean squared error (MSE, lower is better) and peak signal-to-noise ratio (PSNR, in dB, higher is better) for the data sets from Table˜1. PSNR $\infty$ indicates lossless compression.
We first evaluate the compression quality using similarity metrics comparing the compressed representation with their dense texture counterparts. These results are summarized in Table˜4; we tabulate mean squared error (MSE) and peak signal-to-noise-ratio (PSNR) for the test data sets. As expected, the measured quality is a function of sparseness.
@@
While giving a first impression, similarity metrics traditionally have limitations making it hard to draw conclusions relating the metric to the perceived quality. What distinguishes our compression method from other methods is that we encode some values—those that are closer to the VDB background value; i.e., some local features will be perceived as blocky or have missing features, while other volumetric regions will appear as if compressed with a lossless algorithm (cf. Fig.˜2). This is particularly hard to assess from just looking at similarity metrics like MSE or PSNR. What we can conclude from those results though is at which compression rate our algorithm achieves lossless encoding. We observe that many of our data sets can be compressed by 25 % and higher without any quality loss at all.
@@
We then evaluate the impact of the closest, farthest, and median point-in-range metrics from Section˜3.1. We exemplarily present results for the WDAS cloud in Fig.˜3, which indicate superiority of the farthest and median over the closest point-in-range metric. We observed similar outcomes for experiments with the data sets not shown here, leading us to the conclusion that these two metrics seem generally superior.
@@ 5.3 Comparison with Dense Texture Multi-GPU Volume Renderer
We compare the rendering performance of VDBs to that of using random access with dense textures. For that we use the Barney renderer with the extensions described in Section˜4 and either activate or deactivate VDB compression. In the former case, Barney will encode the volume data sets using VDB as described above, while in the latter case the structured-regular grid representation is used. Barney represents structured volumes with 3D CUDA textures.
@@
We present average frame rates with our volume path tracer for the rendered images in Table˜5. These frame rate estimates indicate how long it takes to render noisy convergence frames; the images in Table˜1 represent converged results over 1K convergence frames. To estimate the framerates we chose the $1:8$ compression ratio across all data sets, but note that framerates are stable within small error margins regardless of the size of the VDB. While not faster, we observe that software trilinear interpolation with NanoVDB achieves competitive frame rates compared to using 3D CUDA textures. As the NanoVDBs always fit into one GPU, our algorithm outperforms dense textures on the galaxy data set that requires rendering on three GPUs.
@@ 5.4 Quality Comparison with ZFP
Figure 4: Quality comparison of our method with ZFP. We report mean squared error (MSE, lower is better) and peak signal-to-noise ratio (PSNR, in dB, higher is better) for the data sets from Table˜1, for compression rates between 1-20 %.
@@
We also compare the quality of our compression with that of ZFP. We do not implement a full ZFP renderer due to the technical debt involved. Instead, we compress and then decompress the volume data sets from Table˜1 with ZFP using different compression rates. The decompressed volumes we compare to the original, uncompressed data sets, allowing us to compute MSE and PSNR presented in Fig.˜4.
We are also interested in a qualitative comparison of the two compression algorithms given rendered images. For further analysis we implemented an interval-based implicit ISO surface ray marcher to extract the $0.5$ ISO surface of the WDAS cloud data set. We show renderings with both our algorithm as well as with a ZFP back-and-forth compressed and decompressed structured-regular volume, in Fig.˜5, as well as difference images using F LIP [ANA20].
@@ 5.5 VDB and Dense Texture Back-and-forth Conversion
In Table˜6 we present the results of our experiments. We report the size of the original data converted to NanoVDB and the size of the corresponding structured-regular grid. We further report the smallest size in bytes our algorithm achieves while compressing the data with zero error (MSE=0). For the original resolution, e.g., we observe that compression without error gives a size of 5.8 GB while the original data was 4.1 GB in size. We finally also report the error we realize when compressing the data to the exact size of the original; when compressing the high-resolution data set to match the original size of 4.1 GB, e.g., the MSE we realize is $0.001$.

### v2

# GPU Volume Rendering with VDB Compression

@@ 1 Introduction
In this paper we propose to hierarchically encode—and by doing so compress—the volume data, allowing for per-thread decompression in a GPU compute kernel. We use the sparse voxel representation (VDB) for that. VDBs, so named because they represent volumetric grids that share similarities with B+trees [Mus13], are popular in production rendering and industry-strength libraries exist that are optimized for construction and sampling on NVIDIA GPUs. As such, VDBs are often used for data sets like the one shown in GPU Volume Rendering with VDB Compression.
@@ 3.1 Fixed-Rate Compression Algorithm
Algorithm 1 provides a high-level overview. The input consists of the volume itself, its spatial extent represented by $W$, $H$, and $D$ (in voxels), and the user-provided quality value. We define point sampling routines on the volume itself—without loss of generality, we only sample the volume at exact (integer) voxel positions though (volume.value); as well as routines to activate and set individual voxels on the VDB (Activate).
1:function Compress($volume$, $W$, $H$, $D$, $quality$)
@@
3:     $I=$ argmax($hist$)
4:     $B=volume.valueAt(I)$triangleright background value
@@
6:     $brickSize=$ $int3$(252**52 ∗ ∗ 5)
7:     $numBricks=$ $int3$(${W,H,D}/brickSize$)
@@
11:         $lo=brickID*brickSize$
12:         $hi=(brickID+1)*brickSize$
@@
16:     SortBy($brickRefs$, $valueRanges$, SimilarityFunc($B$))
@@
19:     $bricksToActivate=numBricks*quality$
@@
22:              Activate($Voxel$, $volume.value(Voxel)$)
@@
In the second phase we form bricks of size $32\times 32\times 32$ to cover the original volume, so that each brick covers $64$ leaf nodes of the VDB. We compute value ranges for each brick using the original volume.
@@
$f1$ and $f2$ compute the closest and the farthest distance from the scalar to the range given by $[lo:hi]$, respectively, while $f3$ computes the distance to the median of the range. We refer to the metrics as the closest, farthest and median point-in-range metrics and evaluate their impact on compression in Section 5.
@@ 3.2.1 Compression
We use OpenVDB to implement the fixed-rate compression algorithm. We start from an empty openvdb::FloatGrid. In OpenVDB, grids associate trees, the internal representation of the VDB in memory, with transforms (e.g., voxel to object space). This internal representation is a Tree4<float, 5, 4, 3>, i.e., the tree has four levels using the $\{5,4,3\}$ layout as described above. The tree provides a function addTile(level,Coord(),value) to set and activate cells on each tree level.
@@
Regarding compression performance we note a $1:1$ relationship between activated voxels and final output size. We also note that activating a whole tile (addTile(), see above) on a level higher than leaf node level 00 results in OpenVDB just setting all the voxels in that tile to the same value, so there is no additional compression to gain from using hierarchical or level-of-detail encoding here.
@@ 4 ANARI Renderer
Given our conversion algorithm plus texture sampling routines on the GPU we can implement a renderer. We chose to give the renderer an ANARI API. ANARI support has recently been added to ParaView [WZA24] so our renderer can be evaluated using standard visualization tools and be compared with other renderers that also use the ANARI standard (cf. Fig. 1).
@@ 5 Evaluation
We evaluate the quality and performance of our compressed representation for direct volume rendering. For that we use the structured-regular volume data sets from Table 1. The aneurism data set is particularly sparse. The WDAS cloud data set by Walt Disney Animation Studios [Wal18] is originally available as a VDB and was resampled to a structured-regular representation retaining the original 2K resolution. As we also have access to the original VDB, this gives us the possibility to compare back-and-forth conversion using our algorithm. The airplane data set was simulated with OpenFOAM; we use the Q-criterion vorticity field which is usually sparse compared to other variables. The data was voxelized using the MESIO library [MvJB22], as detailed in [FMB25]. We have access to two versions of that data set: one that represents the vorticity, and a filtered version of that field where $Q\in[0:500K]$, retaining only 4.4 % of the original voxels. We also use a high-resolution, non-cosmological isolated galaxy data set for the evaluation. We utilize an example from [WS23], designed to resemble a Milky Way-type galaxy. The simulation incorporates magnetohydrodynamics (MHD) along with standard subgrid physics commonly employed in galaxy formation studies, including gas dynamics with a magnetic field. The simulation was produces using ChaNga [JGM08] and subsequently converted to a floating-point volume on a structured-regular grid. To simplify the comparison we assume voxels to be represented with 32-bit floating-point values; we convert aneurism and heptane to that format whose voxels use 8-bit values.
@@ 5.1 Compression Rate
We first evaluate if, by using our fixed-rate encoding, the target compression rate is achieved and compress the data set with different ratios from 1:8 to 1:1 (the latter meaning the targeted reduction in size is zero percent). In the latter case the algorithm is guaranteed to be lossless but will only activate voxels that are not considered empty, so the actual size will be lower than the original one. To assess this we report the actual size in memory of the resulting NanoVDBs in Table 2.
@@
We are also interested in the worst case we have to expect when losslessly compressing a volume that is not sparse. This is an extreme case that we simulate by generating synthetic volume data sets. We generate noise textures of different sizes to fill the structured-regular grids with voxels so that no local neighborhood is considered empty by our algorithm. We report results for this experiment in Table 3. For large-scale data we observe that the overhead does not exceed 10 % of the original size. We also found, although without a formal evaluation, that there is no correlation between sparseness and render performance and that NanoVDB performs equally well on data that is dense. This suggests that VDBs are a viable replacement even for dense textures.
@@ 5.2 Compression Quality
Table 4: Compression rate statistics. We report mean squared error (MSE, lower is better) and peak signal-to-noise ratio (PSNR, in dB, higher is better) for the data sets from Table 1. PSNR =absent=$\infty$= ∞ indicates lossless compression.
We first evaluate the compression quality using similarity metrics comparing the compressed representation with their dense texture counterparts. These results are summarized in Table 4; we tabulate mean squared error (MSE) and peak signal-to-noise-ratio (PSNR) for the test data sets. As expected, the measured quality is a function of sparseness.
@@
While giving a first impression, similarity metrics traditionally have limitations making it hard to draw conclusions relating the metric to the perceived quality. What distinguishes our compression method from other methods is that we encode some values—those that are closer to the VDB background value; i.e., some local features will be perceived as blocky or have missing features, while other volumetric regions will appear as if compressed with a lossless algorithm (cf. Fig. 2). This is particularly hard to assess from just looking at similarity metrics like MSE or PSNR. What we can conclude from those results though is at which compression rate our algorithm achieves lossless encoding. We observe that many of our data sets can be compressed by 25 % and higher without any quality loss at all.
@@
We then evaluate the impact of the closest, farthest, and median point-in-range metrics from Section 3.1. We exemplarily present results for the WDAS cloud in Fig. 3, which indicate superiority of the farthest and median over the closest point-in-range metric. We observed similar outcomes for experiments with the data sets not shown here, leading us to the conclusion that these two metrics seem generally superior.
@@ 5.3 Comparison with Dense Texture Multi-GPU Volume Renderer
We compare the rendering performance of VDBs to that of using random access with dense textures. For that we use the Barney renderer with the extensions described in Section 4 and either activate or deactivate VDB compression. In the former case, Barney will encode the volume data sets using VDB as described above, while in the latter case the structured-regular grid representation is used. Barney represents structured volumes with 3D CUDA textures.
@@
We present average frame rates with our volume path tracer for the rendered images in Table 5. These frame rate estimates indicate how long it takes to render noisy convergence frames; the images in Table 1 represent converged results over 1K convergence frames. To estimate the framerates we chose the $1:8$ compression ratio across all data sets, but note that framerates are stable within small error margins regardless of the size of the VDB. While not faster, we observe that software trilinear interpolation with NanoVDB achieves competitive frame rates compared to using 3D CUDA textures. As the NanoVDBs always fit into one GPU, our algorithm outperforms dense textures on the galaxy data set that requires rendering on three GPUs.
@@ 5.4 Quality Comparison with ZFP
Figure 4: Quality comparison of our method with ZFP. We report mean squared error (MSE, lower is better) and peak signal-to-noise ratio (PSNR, in dB, higher is better) for the data sets from Table 1, for compression rates between 1-20 %.
@@
We also compare the quality of our compression with that of ZFP. We do not implement a full ZFP renderer due to the technical debt involved. Instead, we compress and then decompress the volume data sets from Table 1 with ZFP using different compression rates. The decompressed volumes we compare to the original, uncompressed data sets, allowing us to compute MSE and PSNR presented in Fig. 4.
We are also interested in a qualitative comparison of the two compression algorithms given rendered images. For further analysis we implemented an interval-based implicit ISO surface ray marcher to extract the $0.5$ ISO surface of the WDAS cloud data set. We show renderings with both our algorithm as well as with a ZFP back-and-forth compressed and decompressed structured-regular volume, in Fig. 5, as well as difference images using F LIP [ANA20].
@@ 5.5 VDB and Dense Texture Back-and-forth Conversion
In Table 6 we present the results of our experiments. We report the size of the original data converted to NanoVDB and the size of the corresponding structured-regular grid. We further report the smallest size in bytes our algorithm achieves while compressing the data with zero error (MSE=0). For the original resolution, e.g., we observe that compression without error gives a size of 5.8 GB while the original data was 4.1 GB in size. We finally also report the error we realize when compressing the data to the exact size of the original; when compressing the high-resolution data set to match the original size of 4.1 GB, e.g., the MSE we realize is $0.001$.